# Mind the Gap: Navigating Inference with Optimal Transport Maps

### v1

@@ 4.1 Constructing the classifier
The identification of jet origin, commonly referred to as jet tagging, is a key challenge in particle physics, where machine learning techniques have become increasingly prominent. In this work, we construct a classifier based on a transformer [transformer] architecture that closely resembles the state-of-the-art models currently employed for jet-tagging at LHC experiments [GN2, UParT].
The classifier is designed to estimate the conditional probability distribution for the origin of a jet, $p(l\mid\{\vec{c}_{\text{obs}}\},\text{jet}_{\text{obs}})$, where $\vec{c}_{\text{obs}}$ denotes the set of features associated with the jet constituents, and $\text{jet}_{\text{obs}}$ is a fixed-size vector of jet-level observables, including the pseudorapidity $\eta$, azimuthal angle $\phi$, transverse momentum $p_{T}$, energy $E$, and constituent multiplicity $N$. A complete list of input variables, along with their corresponding distributions, is provided in Appendix C.
The classifier uses a dedicated ClassToken [cls_token] to aggregate the input set $\vec{c}_{\text{obs}}$ into a fixed 512-dimensional latent representation. The Transformer encoder is composed of multiple stacked layers, each consisting of a multi-head self-attention mechanism followed by a feed-forward neural network [transformer]. The jet input features $\text{jet}_{\text{obs}}$ are integrated into the encoder via Feature-wise Linear Modulation (FiLM), following the methodology established in Ref. [FiLM_newer].
The jet’s latent representation is passed to the output multi-layer perceptron head (MLP Head), which consists of a sequence of linear layers with GELU activation functions [gelu], each maintaining a fixed latent dimensionality of 128. The final output of the network is a 10-dimensional latent vector containing the unnormalized class scores (logits) corresponding to the different jet categories. An overview of the classifier architecture is shown in Fig. 2. The first 128-dimensional layer of the MLP Head, denoted $z_{128}$, is the target of the calibration. The final 128-dimensional layer ($z^{\prime}_{128}$) and the output layer ($z^{\prime}_{10}$) are also shown in Fig. 2 with dashed lines, and are used in the evaluation of the calibration procedure, as further discussed in Section 5.
Figure 2: The two diagrams show (a) the transformer encoder and (b) the output multi-layer perceptron head (MLP Head). The transformer architecture incorporates a sequence of encoder layers [prelayer_norm], each comprising multi-head self-attention mechanisms (MHA) [transformer] followed by feed-forward neural networks (MLP) [shazeer2020gluvariantsimprovetransformer] with Feature-wise Linear Modulation (FiLM) [FiLM, FiLM_newer] for conditional integration. The MLP Head consists of a sequence of linear transformations and GELU [gelu] with a final output size of $10$ for the class probabilities.
The model is trained using the publicly available JetClass dataset [JetClass]. Details of the training procedure are provided in Appendix B.
@@ 4.2 Deriving the optimal transport map
The primary objective of this work is to derive an optimal transport map, denoted by $\widehat{T}_{\omega}$, such that
@@
where $p_{\text{sim}}$ and $p_{\text{data}}$ represent the distributions of the latent representation $z_{128}$ in simulation and real data, respectively, conditioned on $\omega$. The conditioning variable $\omega$ typically includes jet kinematic features such as jet-$p_{T}$, jet-$\eta$, and other relevant labels extracted from the simulated dataset. In this work we neglect dependencies on the jet kinematics and thus restrict $\omega$ to the $q/g$ label of the simulated sample that we aim to calibrate.
The optimal transport map $\widehat{T}_{\omega}$ is a unique transformation that minimally distorts $p_{\text{sim}}$ in order to match $p_{\text{data}}$. Refs. [Original_ot_solution, cpflows, Algren_2024, ATLAS-PAPER-2025, DBLP:journals/corr/abs-2106-01954, bunne2023supervisedtrainingconditionalmonge] have shown that $\widehat{T}_{\omega}$ can be learned by minimizing the Euclidean transport cost between the simulated and observed data distributions, using a pair of input-convex neural networks (ICNNs) [ICNN]. This approach is formalized by the following optimization problem:
@@
Where $f$ and $g$ are two distinct input-convex neural networks (ICNNs), and $z\sim p_{\text{sim}}$ and $y\sim p_{\text{data}}$ denote samples from the simulated and target distributions, respectively. For simplicity, we have omitted the subscript 128128{128}128 from $z$ and $y$.
At convergence, the optimal transport map is given by $\widehat{T}_{\omega}(z|\omega)=\nabla_{z}g(z|\omega)$, which maps simulated latent representations $z$ to their calibrated counterparts $y$, conditioned on $\omega$. Further details on the ICNN architecture and training procedure are provided in Appendices A and B.
@@ 5.1 Evaluating closure in the latent space
To visualize the impact of calibration within the high-dimensional latent space $z_{128}$, Principal Component Analysis (PCA) [PCA] is applied to reduce the dimensionality to the five leading principal components. The distributions of these components are presented as corner plots in Fig. 3.
Fig. 3(a) illustrates the presence of the expected discrepancies between the source (simulated) and target (synthetic) data distributions. After applying the optimal transport map, Fig. 3(b) shows a significant reduction in these discrepancies, indicating successful alignment between the latent representations of the two domains.
@@
The calibration procedure is further evaluated using discriminators, denoted $\mathcal{D}_{L}$, trained to distinguish between the source and target distributions across the initial 128-dimensional layer ($z_{128}$), the final 128-dimensional layer ($z^{\prime}_{128}$), and the output layer ($z^{\prime}_{10}$) of the MLP Head. Fig. 4(a) presents the receiver operating characteristic (ROC) curves of $\mathcal{D}_{L}$ trained on each of these layers. As expected, discrepancies between the source and target distributions are largest in $z_{128}$ and decrease progressively in deeper layers, with the smallest differences observed in $z^{\prime}_{10}$.
@@
The discriminant score produced by $\mathcal{D}_{L}$ in the $z_{128}$ latent space is presented in Fig. 4(b). The source and target distributions exhibit only limited overlapping support and are otherwise nearly disjoint. In such cases, calibration methods based on reweighting—referred to as vertical calibration in the introduction are subject to substantial statistical inefficiency due to the sparse overlap, thereby limiting their applicability. By contrast, the proposed horizontal calibration approach is not hindered by the lack of distributional overlap and thus remains a feasible strategy to calibrate high dimensional latent representations.
@@
To validate this, the impact of the calibration is further examined in the classifier’s output space. In typical applications output probabilities are combined to construct log-likelihood discriminants. Fig. 5 displays the distributions of these discriminants, where Higgs boson decays to $b$- or $c$-quark pairs are considered signal, while multijet and top-quark events define the background hypothesis.
Prior to calibration, clear discrepancies is observed between the source (blue) and target (black) distributions. As expected, the application of the calibration procedure (red) substantially reduces these differences, indicating improved agreement between the two distributions.
[…]
ıin p_Hbb_p_QCD,p_Hcc_p_QCD,p_Hcc_p_Tbqq
@@ Acknowledgements
Appendix A Optimal transport
@@
The classical primal formulation of the optimal transport problem [monge1781memoire], is defined as:
@@
[…]
@@
where $f$ and $g$ are functions subject to the constraint $f(y)+g(z)\leq c(y,z)$ [villani2008optimal]. Ref. [Original_ot_solution] demonstrated that for continuous distributions, the optimal transport can be derived through the following optimization problem:
@@
Figure 6: The diagram shows the architecture of the ICNN[ICNN] used to derive the optimal transport by optimizing Eq. 10. The specific implementation of the ICNN that is used in this paper follows the architecture described in [cpflows, Algren_2024].
A.1 Method to derive the optimal transport
As established in Appendix A, the solution to Eq. 10 requires two convex functions $f$ and $g$. While conventional neural networks (NN) are not limited to convex functions, Ref. [ICNN] introduced input-convex neural networks (ICNNs), which limit the parametrization of NNs to only convex functions $f(y;\theta)$ with respect to their input variables $y$. The ICNN architecture implements a recursive formulation across layers $i=0,...,k-1$, wherein the $i^{\text{th}}$ layer is defined as:
@@ A.1 Method to derive the optimal transport
where $h_{0}=W^{h}_{i}=0$, and $\sigma$ represents a convex, monotonically non-decreasing activation function. A schematic representation of the ICNN architecture is illustrated in Fig. 6.
There is an extensive literature on the ICNN and using it to derive the optimal transport [Original_ot_solution, cpflows, Algren_2024, ATLAS-PAPER-2025, DBLP:journals/corr/abs-2106-01954, bunne2023supervisedtrainingconditionalmonge] with different tweaks to the ICNN architecture and the training procedure.
The specific architecture and training implementation of the ICNNs used in this paper follow the architecture of Ref. [cpflows] and training procedure of Ref. [Algren_2024].
Appendix B Hyperparameters
B.1 Classifier trained on the JetClass dataset
Transformer based classifier trained on JetClass dataset. The output of the classifier is a 10-dimensional vector representing the probability of the jet belonging to one of the 10 classes. The classifier was trained with 100.000 optimizer steps of AdamW [adamw] with a learning rate of $10^{-4}$ and a warmup of 10.000 steps. The full hyperparameters of the classifier are listed in Table 1.
@@ B.1 Classifier trained on the JetClass dataset
Table 1: Hyperparameters of the classifier trained on JetClass. Diagram of the architecture can be seen in Fig. 2.
B.2 Hyperparameters of the ICNN
Two ICNN networks are used to derive the optimal transport map, both with the same hyperparameters, which are listed in Table 2. The ICNNs are trained on source and target samples of size 8,181,000, with a validation size of 909,000 for each source and target. The latent space is standardized to have zero mean and unit variance before being passed to the ICNNs.
@@ B.2 Hyperparameters of the ICNN
B.3 Hyperparameters of the classifiers used for evaluation $\mathcal{D}_{L}$
In the evaluation stage of the latent space calibration, a classifier is trained to The classifier $\mathcal{D}_{L}$ is the classifier used to distinguish between the target and source distributions and calibrated and target. The hyperparameters of $\mathcal{D}_{L}$ are listed in Table 3. The total number of parameters in $\mathcal{D}_{L}$ is around a $1$ million.

### v2

@@ 5.1 Constructing the classifier
The identification of jet origin, commonly referred to as jet tagging, is a key challenge in particle physics, where machine learning techniques have become increasingly prominent. In this work, we construct a classifier based on a transformer [56] architecture that closely resembles the state-of-the-art models currently employed for jet-tagging at LHC experiments [30, 29].
The classifier is designed to estimate the conditional probability distribution for the origin of a jet, $p(l\mid\{\vec{c}_{\text{obs}}\},\text{jet}_{\text{obs}})$, where $\vec{c}_{\text{obs}}$ denotes the set of features associated with the jet constituents, and $\text{jet}_{\text{obs}}$ is a fixed-size vector of jet-level observables, including the pseudorapidity $\eta$, azimuthal angle $\phi$, transverse momentum $p_{T}$, energy $E$, and constituent multiplicity $N$. A complete list of input variables, along with their corresponding distributions, is provided in Appendix E.
The classifier uses a dedicated ClassToken [57] to aggregate the input set $\vec{c}_{\text{obs}}$ into a fixed 512-dimensional latent representation. The Transformer encoder is composed of multiple stacked layers, each consisting of a multi-head self-attention mechanism followed by a feed-forward neural network [56]. The jet input features $\text{jet}_{\text{obs}}$ are integrated into the encoder via Feature-wise Linear Modulation (FiLM), following the methodology established in Ref. [58].
The jet’s latent representation is passed to the output multi-layer perceptron head (MLP Head), which consists of a sequence of linear layers with GELU activation functions [59], each maintaining a fixed latent dimensionality of 128. The final output of the network is a 10-dimensional latent vector containing the unnormalized class scores (logits) corresponding to the different jet categories. An overview of the classifier architecture is shown in Fig.˜2. The first 128-dimensional layer of the MLP Head, denoted $z_{128}$, is the target of the calibration. The final 128-dimensional layer ($z^{\prime}_{128}$) and the output layer ($z^{\prime}_{10}$) are also shown in Fig.˜2 with dashed lines, and are used in the evaluation of the calibration procedure, as further discussed in Section 6.
Figure 2: The two diagrams show (a) the transformer encoder and (b) the output multi-layer perceptron head (MLP Head). The transformer architecture incorporates a sequence of encoder layers [60], each comprising multi-head self-attention mechanisms (MHA) [56] followed by feed-forward neural networks (MLP) [61] with Feature-wise Linear Modulation (FiLM) [62, 58] for conditional integration. The MLP Head consists of a sequence of linear transformations and GELU [59] with a final output size of 1010 for the class probabilities.
The model is trained using the publicly available JetClass dataset [50]. Details of the training procedure are provided in Appendix D.
@@ 5.2 Deriving the optimal transport map
The primary objective of this work is to derive an optimal transport map, denoted by $\widehat{T}_{\omega}$ [63, 9], such that
@@
where $p_{\text{sim}}$ and $p_{\text{data}}$ represent the distributions of the latent representation $z_{128}$ in simulation and real data, respectively, conditioned on $\omega$. The conditioning variable $\omega$ typically includes jet kinematic features such as $p_{T}, jet$-$\eta$, and other relevant labels extracted from the simulated dataset [12]. In this work we neglect dependencies on the jet kinematics and thus restrict $\omega$ to the $q/g$ label of the simulated sample that we aim to calibrate.
The optimal transport map $\widehat{T}_{\omega}$ is a unique transformation that minimally distorts $p_{\text{sim}}$ in order to match $p_{\text{data}}$ [63, 9, 10]. Refs. [15, 16, 46, 12, 17, 18] have shown that $\widehat{T}_{\omega}$ can be learned by minimizing the Euclidean transport cost between the simulated and observed data distributions, using a pair of input-convex neural networks (ICNNs) [64, 16]. This approach is formalized by the following optimization problem[10]:
@@
Where $f$ and $g$ are two distinct input-convex neural networks (ICNNs), $z\sim p_{\text{sim}}$ and $y\sim p_{\text{data}}$ denote samples from the simulated and target distributions, respectively. For simplicity, we have omitted the subscript 128{128} from $z$ and $y$.
At convergence, the optimal transport map is given by $\widehat{T}_{\omega}(z|\omega)=\nabla_{z}g(z|\omega)$, which maps simulated latent representations $z$ to their calibrated counterparts $y$, conditioned on $\omega$ [18, 12, 46].
Further details on the ICNN architecture and training procedure are provided in Appendices B and D.
@@ 6.1 Evaluating closure in the latent space
To visualize the impact of calibration within the high-dimensional latent space $z_{128}$, Principal Component Analysis (PCA) [65] is applied to reduce the dimensionality to the five leading principal components. The distributions of these components are presented as corner plots in Fig.˜3.
Fig.˜3(a) illustrates the presence of the expected discrepancies between the source (simulated) and target (synthetic) data distributions. After applying the optimal transport map, Fig.˜3(b) shows a significant reduction in these discrepancies, indicating successful alignment between the latent representations of the two domains.
@@
The calibration procedure is further evaluated using discriminators, denoted $\mathcal{D}_{L}$, trained to distinguish between the source and target distributions across the initial 128-dimensional layer ($z_{128}$), the final 128-dimensional layer ($z^{\prime}_{128}$), and the output layer ($z^{\prime}_{10}$) of the MLP Head. Fig.˜4(a) presents the receiver operating characteristic (ROC) curves of $\mathcal{D}_{L}$ trained on each of these layers. Discrepancies between the source and target distributions are largest in $z_{128}$ and decrease progressively in deeper layers, with the smallest differences observed in $z^{\prime}_{10}$.
@@
The discriminant score produced by $\mathcal{D}_{L}$ in the $z_{128}$ latent space is presented in Fig.˜4(b). The source and target distributions exhibit only limited overlapping support and are otherwise nearly disjoint. In such cases, calibration methods based on reweighting—referred to as vertical calibration in section Section˜2.1 are subject to substantial statistical inefficiency due to the sparse overlap, thereby limiting their applicability. By contrast, the proposed horizontal calibration approach is not hindered by the lack of distributional overlap and thus remains a feasible strategy to calibrate high dimensional latent representations.
@@
To validate this, the impact of the calibration is further examined in the classifier’s output space. In typical applications output probabilities are combined to construct log-likelihood discriminants. Fig.˜5 displays the distributions of these discriminants, where Higgs boson decays to $b$- or $c$-quark pairs are considered signal, while multijet and top-quark events define the background hypothesis.
Prior to calibration, clear discrepancies is observed between the source (blue) and target (black) distributions. The application of the calibration procedure (red) substantially reduces these differences, indicating improved agreement between the two distributions.
@@ Acknowledgements
Appendix A Reweighting dilution factors
A.1 General dilution factors
The dilution factor $\rho$ (such that for $N$ samples from $p_{1}$, the effective number of reweighted simulated samples is $N_{\mathrm{eff}}\equiv N_{\mathrm{sim}}/\rho$) for $x\sim p_{1}$ reweighted to $p_{2}$ via the ratio $s(x)\equiv p_{2}(x)/p_{1}(x)$ is given by
$\frac{\mathbb{E}_{x\sim p_{1}}[s(x)^{2}]}{\mathbb{E}_{x\sim p_{1}}[s(x)]^{2}}.$
The denominator,
$\left(\int_{x}p_{1}(x)\frac{p_{2}(x)}{p_{1}(x)}\right)^{2},$
is unity, since $p_{2}(x)$ is a probability density. The numerator is given by
$\int_{x}p_{1}(x)\left(\frac{p_{2}(x)}{p_{1}(x)}\right)^{2}=\int_{x}p_{2}(x)\frac{p_{2}(x)}{p_{1}(x)},$
i.e. the expected value over $p_{2}$ of $s$. To summarize, the dilution factor is
$\rho=\mathbb{E}_{x\sim p_{2}}[s(x)]$
and is related to the $\chi^{2}$-divergence [21, 22]: $\mathbb{E}_{x\sim p_{2}}[s(x)]=1+\chi^{2}(p_{2}||p_{1})$. The minimum $\rho$ is unity when $\forall x.~p_{1}(x)=p_{2}(x)$, i.e. no reweighting is required.
A.2 Dilution factors for multivariate normal distributions
To make the previous discussion concrete, we derive analytic expressions for $\rho$ for two multivariate normal distributions (MVNs). The log-density ratio of two MVNs $p_{1}$ and $p_{2}$, with respective means $\mu_{1},\mu_{2}$ and covariances $\Sigma_{1},\Sigma_{2}$, is given by
$2\log\frac{p_{2}(x)}{p_{1}(x)}=\log\frac{|\Sigma_{1}|}{|\Sigma_{2}|}-(x-\mu_{2})^{T}\Sigma_{2}^{-1}(x-\mu_{2})+(x-\mu_{1})^{T}\Sigma_{1}^{-1}(x-\mu_{1}).$
$\mathbb{E}_{x\sim p_{2}}[s(x)]$ in this case is the result of a Gaussian integral. To ease notation, we define
$A\equiv 2\Sigma_{2}^{-1}-\Sigma_{1}^{-1}\qquad b\equiv 2\Sigma_{2}^{-1}\mu_{2}-\Sigma_{1}^{-1}\mu_{1}\qquad c\equiv\mu_{2}^{T}\Sigma_{2}^{-1}\mu_{2}-\frac{1}{2}\mu_{1}^{T}\Sigma_{1}^{-1}\mu_{1},$
yielding
$\rho$ $=\frac{\sqrt{|\Sigma_{1}|}}{|\Sigma_{2}|}\int_{\mathbb{R}^{n}}\exp\left(-\frac{1}{2}x^{T}Ax+b^{T}x-c\right)$
$=\frac{\sqrt{|\Sigma_{1}|}}{|\Sigma_{2}|\sqrt{|A|}}\exp\left(\frac{1}{2}b^{T}A^{-1}b-c\right).$
$\rho$ is only finite if $A$ is positive definite. In case $\Sigma_{1}$ and $\Sigma_{2}$ are simultaneously diagonal, then the dilution factor is infinite unless $\forall i.~A_{ii}>0$, i.e. in every dimension the variance of $p_{1}$ can be no less than half that of $p_{2}$. This prevents the vertical calibration procedure (and reweighting in general) from being viable in a wide variety of common scenarios.
If $\Sigma_{1}=\Sigma_{2}=\Sigma$, then $\rho$ is simply $\exp(\Delta^{T}\Sigma\Delta)$ with $\Delta\equiv\mu_{1}-\mu_{2}$. In this case the dilution factor grows exponentially with the Mahalanobis distance between the means [23]. If $\mu_{1}=\mu_{2}=\mu$, then
$\rho=\frac{1}{|\Sigma_{2}|}\sqrt{\frac{|\Sigma_{1}|}{|A|}};$
this depends only on the relative covariances and, as stated earlier, diverges when $A$ is not positive definite.
Appendix B Optimal transport with input-convex neural networks
@@
The classical primal formulation of the optimal transport problem [9], is defined as:
@@
where $T$ is the transport map and $\|\cdot\|$ is the cost function that ranks the maps. It is an optimization problem that seeks to find the map $T$ that minimizes $\|\cdot\|$ over all maps that satisfy $T_{\#}p_{\text{sim}}=p_{\text{data}}$, where $T_{\#}$ is the pushforward operator [10]. Monge’s transport maps are defined as deterministic maps, which imposes significant analytical challenges [63, 10]. Consequently, Kantorovich proposed a relaxation of Monge’s primal formulation by conceptualizing transport as a probability measure rather than a deterministic mapping [63]. Additionally to the relaxation, Kantorovich introduced a dual problem that has the same solution as the primal one. The dual problem is defined as:
@@
where $f$ and $g$ are functions subject to the constraint $f(y)+g(z)\leq c(y,z)$ [10]. Ref. [15] demonstrated that for continuous distributions, the optimal transport can be derived through the following optimization problem:
@@
Figure 6: The diagram shows the architecture of the ICNN[64] used to derive the optimal transport by optimizing Eq.˜10. The specific implementation of the ICNN that is used in this paper follows the architecture described in [16, 46].
B.1 Method to derive the optimal transport
As established in Appendix˜B, the solution to Eq.˜10 requires two convex functions $f$ and $g$. While conventional neural networks (NN) are not limited to convex functions, Ref. [64] introduced input-convex neural networks (ICNNs), which limit the parametrization of NNs to only convex functions $f(y;\theta)$ with respect to their input variables $y$. The ICNN architecture implements a recursive formulation across layers i=0,,k1i=0,...,k-1, wherein the $i^{\text{th}}$ layer is defined as:
@@ B.1 Method to derive the optimal transport
where $h_{0}=W^{h}_{i}=0$, and $\sigma$ represents a convex, monotonically non-decreasing activation function. A schematic representation of the ICNN architecture is illustrated in Fig.˜6.
There is an extensive literature on the ICNN and using it to derive the optimal transport [15, 16, 46, 12, 17, 18] with different tweaks to the ICNN architecture and the training procedure.
The specific architecture and training implementation of the ICNNs used in this paper follow the architecture of Ref. [16] and training procedure of Ref. [46].
Figs.˜7 and 8 show the marginal distributions of the input features normalized with $QuantileTransformer$.
Appendix C Results
C.1 Calibrated results for MLP Head
Fig.˜9 show the ROC curves of the discriminators trained to discriminate between the calibration and target distributions, as well as the nominal and target distributions.
Fig.˜10 show the marginal distributions of the output space of the classifier, comparing the target distribution, the transported distribution, and the source distribution. Fig.˜11 shows an additional physics motivated discriminator.
Figure 11: Comparison between the marginal distributions of a physics-motivated one-dimensional discriminant between hadronic top-quark decays and QCD multijet. The distributions are represented as follows: target distribution (black), calibrated distribution (red), and source distribution (blue). The calibration is derived on the $z_{128}$ latent space, then the subsequent MLP Head layers are applied to map it to output space $z_{10}^{{}^{\prime}}$ where Softmax is applied to map the output space to a simplex. Afterwards the output space is projected to physics-motivated one-dimensional discriminants.
C.2 Corner plots in the output space
Figs.˜12 and 13 show the corner plots of the output space of the classifier with and without calibration.
Appendix D Hyperparameters
D.1 Classifier trained on the JetClass dataset
Transformer based classifier trained on JetClass dataset. The output of the classifier is a 10-dimensional vector representing the probability of the jet belonging to one of the 10 classes. The classifier was trained with 100.000 optimizer steps of AdamW [66] with a learning rate of $10^{-4}$ and a warmup of 10.000 steps. The full hyperparameters of the classifier are listed in Table˜1.
@@ D.1 Classifier trained on the JetClass dataset
Table 1: Hyperparameters of the classifier trained on JetClass. Diagram of the architecture can be seen in Fig.˜2.
D.2 Hyperparameters of the ICNN
Two ICNN networks are used to derive the optimal transport map, both with the same hyperparameters, which are listed in Table˜2. The ICNNs are trained on source and target samples of size 8,181,000, with a validation size of 909,000 for each source and target. The latent space is standardized to have zero mean and unit variance before being passed to the ICNNs.
@@ D.2 Hyperparameters of the ICNN
D.3 Hyperparameters of the classifiers used for evaluation $\mathcal{D}_{L}$
In the evaluation stage of the latent space calibration, a classifier is trained to The classifier $\mathcal{D}_{L}$ is the classifier used to distinguish between the target and source distributions and calibrated and target. The hyperparameters of $\mathcal{D}_{L}$ are listed in Table˜3. The total number of parameters in $\mathcal{D}_{L}$ is around a $1$ million.